Find the 7th term of the sequence whose nth term is given by the formula tn = n(n + 2)
Answer by Edwin McCravy(20081)

tn = n(n + 2)

Just substitute 7 for n and simplify

t7 = 7(7 + 2) = 7(9) = 63

Edwin